# Towards Robust and Realible Multimodal Fake News Detection with Incomplete Modality

### v1

@@ 3.3.1. Expert of Text-Modality
(3) $F_{T}=f_{A}^{T}(h_{\text{{[CLS]}}})$
(4) ${y}^{h}=\text{softmax}(\boldsymbol{W}F_{T}+\boldsymbol{b})$
@@
3.3.2. Expert of Image-Modality
@@
(6) $F_{I}=f_{A}^{I}({r}_{\text{{[CLS]}}})$
(7) ${y}^{r}=\text{softmax}(\boldsymbol{W}F_{I}+\boldsymbol{b})$
@@
Figure 3. Distribution of features in space for complete and incomplete modality.:
3.3.3. Expert of Multi-Modality
Modeling the distribution between text modality and image modality is advantageous for capturing deep-level deceptive information. However, incomplete modality information can mislead the modeling of inter-modality distributions, thereby hindering the detection of deceptive information. In this work, we employ a transformer encoder to model the joint expert distribution across multiple modalities.
@@ 3.3.3. Expert of Multi-Modality
(8) $F=(H,R)=(h_{\text{[CLS]}},h_{1},\dots,h_{m},r_{\text{[CLS]}},r_{1},\dots,r_{n})$
@@
(9) $\hat{F}=(\hat{h}_{\text{[CLS]}},\hat{h}_{1},\dots,\hat{h}_{m},\hat{r}_{\text{[CLS]}},\hat{r}_{1},\dots,\hat{r}_{n})=\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d_{k}}}\right)V$
@@
(10) $p_{t},p_{v}=\text{softmax}(\boldsymbol{W}(\hat{h}_{\text{[CLS]}},\hat{r}_{\text{[CLS]}})+b)$
(11) $f=p_{t}\hat{h}_{\text{[CLS]}}+p_{v}\hat{r}_{\text{[CLS]}})$
(12) $y^{f}=\text{softmax}(\boldsymbol{W}f+\boldsymbol{b})$
@@
3.3.4. Multi-Expert Reasoning Paradigm
@@
(13) ${y}^{o}=\sum_{M\in\left\{h,r,f\right\}}{\lambda}_{o}^{M}{y}^{M},{\lambda}_{o}\in\left[0,1\right]$
@@
3.4. Incomplete Modality Adapters
Mainstream MFND tasks typically employ models pre-trained on large-scale datasets as encoders, which may lead to overfitting of feature distributions, a problem that becomes more pronounced when modality information is missing. Inspired by (Gao et al., 2023), we introduce lightweight feature adapters for single-modality image and text, enabling the learning of new modality features to compensate for the missing information while preserving the original distributions learned by the pre-trained models.
@@ 3.4. Incomplete Modality Adapters
(14) $F_{I}=\alpha A^{I}({r}_{\text{{[CLS]}}})+(1-\alpha){r}_{\text{{[CLS]}}}=f_{A}^{I}({r}_{\text{{[CLS]}}})$
(15) $F_{T}=\alpha A^{T}({h}_{\text{{[CLS]}}})+(1-\alpha){h}_{\text{{[CLS]}}}=f_{A}^{T}({h}_{\text{{[CLS]}}})$
@@
3.5. Modality Missing Learning
@@
To enhance the performance of MMLNet in MFND tasks under conditions of incomplete modality information, we propose a modality missing learning algorithm based on a vanilla multimodal contrastive learning framework. This simple yet effective approach aims to mitigate the negative effects of missing modality information by constraining the feature distributions across multiple modalities. Specifically, for each training sample $x$, we first define the corresponding positive sample set $\mathcal{S}_{p}$ and negative sample set $\mathcal{S}_{n}$ for it. Here, $\mathcal{S}_{p}$ contains samples with the same label as $x$, while $\mathcal{S}_{n}$ contains samples with the opposite label as $x$. As shown in Figure 3, we aim to learn a robust latent representation of $x$ by attracting $x$ to the representation of samples in $\mathcal{S}_{p}$ and repidating it away from the representation of samples in $\mathcal{S}_{n}$ as follows:
(16) $\mathcal{L}_{m}=\frac{1}{\left|\mathcal{S}_{p}\right|}\sum_{p\in\mathcal{S}_{p}}-\log\frac{\exp\left(f(\mathbf{h})\cdot f\left(\mathbf{h}_{p}\right)/\tau\right)}{\sum_{n\in\mathcal{S}_{n}}\exp\left(f(\mathbf{h})\cdot f\left(\mathbf{h}_{n}\right)/\tau\right)}$
@@
However, in incomplete modal MFND task, it is suboptimal to weight all positive and negative samples equally. This alignment objective is not fully compatible with multimodal fake news detection with incomplete modalities. Based on the aforementioned perceptions, our model encounters two common but difficult problems caused by the lack of modality information: 1) distinguishing samples with similar semantic information but opposite labels due to incomplete modal information, and 2) narrowing the gap between samples with different semantic information but the same label due to incomplete modal information. To accomplish this, we introduce a label-aware adaptive weighting strategy. Specifically, we measure the semantic similarity of two samples by measuring the cosine similarity, each sample is re-weighted as follows:
(19) $w_{c}=\left\{\begin{array}[]{l}1-cos(h_{c},h),\left(c\in\mathcal{S}_{p}\right)\\ 1+cos(h_{c},h),\left(c\in\mathcal{S}_{n}\right)\end{array}\right$.
@@
(20) $\hat{\mathcal{L}}_{m}=\frac{1}{\left|\mathcal{S}_{p}\right|}\sum_{p\in\mathcal{S}_{p}}-\log\frac{w_{p}\cdot\exp\left(f(\mathbf{h})\cdot f\left(\mathbf{h}_{p}\right)/\tau\right)}{\sum_{n\in\mathcal{S}_{n}}w_{n}\cdot\exp\left(f(\mathbf{h})\cdot f\left(\mathbf{h}_{n}\right)/\tau\right)}$
@@
In fact, modality missing learning is not only effective for multimodal features. We also apply it to text modality features $F_{T}$ and image modality features $F_{I}$ to further constrain the distribution of unimodal features and enhance the robustness of the MMLNet model in the incomplete modality MFND task.
3.6. Overall Optimization
@@
(22) $\mathcal{L}=\sum_{M\in\{I,T,F\}}\lambda_{o}^{M}\left(\lambda_{c}\mathcal{L}_{c}^{M}+\lambda_{m}\mathcal{L}_{m}^{M}\right)$
@@
4. Experiments
In this section, we conduct relevant experiments to address several research questions. Through these experimental results, we systematically evaluate the effectiveness of MMLNet in tackling key challenges in incomplete modality fake news detection:
  • RQ1: Can MMLNet effectively enhance the performance of fake news detection?
  • RQ2: Is the proposed design robust for fake news detection under incomplete modalities?
  • RQ3: Do individual components contribute to improving fake news detection capability?
  • RQ4: Can modallity missing learning significantly improve incomplete modality MFND?
@@ 4. Experiments
4.1. Datasets
We evaluate the performance of MMLNet on three real-world datasets: Weibo (Jin et al., 2017), Weibo21 (Nan et al., 2021), and Pheme (Zubiaga et al., 2017), covering both Chinese and English. The Weibo dataset is the most widely used Chinese fake news detection dataset, containing fake and real news from Sina Weibo between 2012 and 2016, verified by authoritative sources such as Xinhua News Agency and Sina’s official debunking system. The Weibo21 dataset, released in 2021, is an updated version of the Weibo dataset that includes more recent Weibo news. Pheme is an English-language dataset consisting of tweets from Twitter, focusing on five breaking news events. Each event comprises a set of posts containing a substantial amount of textual and visual content, along with corresponding annotations. Three events are randomly selected, and their associated tweets are used for training, while tweets from the remaining events are utilized for testing. For comparison with baseline methods, we follow previous studies (Liu et al., 2025; Dong et al., 2024; Zhang et al., 2024b; Zhao et al., 2023) and use the same preprocessed dataset.
(a) Text ↓0%, Image ↓100%.
(b) Text ↓25%, Image ↓75%.
(c) Text ↓50%, Image ↓50%.
(d) Text ↓75%, Image ↓25%.
(e) Text ↓100%, Image ↓0%.
Figure 4. Experimental results of Weibo dataset.
(a) Text ↓0%, Image ↓100%.
(b) Text ↓25%, Image ↓75%.
(c) Text ↓50%, Image ↓50%.
(d) Text ↓75%, Image ↓25%.
(e) Text ↓100%, Image ↓0%.
Figure 5. Experimental results of Weibo21 dataset.
(a) Text ↓0%, Image ↓100%.
(b) Text ↓25%, Image ↓75%.
(c) Text ↓50%, Image ↓50%.
(d) Text ↓75%, Image ↓25%.
(e) Text ↓100%, Image ↓0%.
Figure 6. Experimental results of Pheme dataset.
4.2. Evaluation Metrics
The multimodal fake news detection problem is a binary classification task. Based on previous work (Dong et al., 2024; Liu et al., 2025, 2024), we apply the accuracy score (ACC), the area under the receiver operating characteristic curve (AUC), and the macro-F1 score (M-F1) as evaluation metrics.
4.3. Baseline
We compare the two open-sourced sota MFDN methods including NSLM (Dong et al., 2024) and MIMoE (Liu et al., 2025). We also fine-tunes the parameters of MLLM to show the ability of MLLM in MFND in complete and incomplete modal scenarios, specifically using the open-sourced MLLM Qwen-VL-7B [^1] https://github.com/QwenLM/Qwen-VL and LLaVA-v1.5-13B [^2] https://huggingface.co/liuhaotian/llava-v1.5-13b.
4.4. Implementation Details
To facilitate future research on incomplete modality MFND, for handling missing text modality, we set a random seed and pre-store the missing text. For missing image modality, we store the indices of the missing patches. Specifically, the image is resized to $224\times 224$ and decompose it into patches of size $32\times 32$. For the simplicity of the experiment, the training set and the test set have the same missing settings.
@@ 4.4. Implementation Details
4.5. Main Results
Answear to RQ1. We conducted experiments using Chinese and English datasets in the complete modality scenario, as shown in Table 1. MMLNet significantly outperforms existing sota MFND methods and MLLM in three evaluation indicators. On the Chinese weibo dataset, the accuracy scores of the sota MFND method and MLLM are improved by 2.73% and 12.22% respectively. On the English pheme dataset, the accuracy scores of the sota MFND method and MLLM are improved by 2.14% and 30.68% respectively. These experimental results prove that MMLNet can more accurately identify misinformation in news. For example, the accuracy scores on Weibo and Pheme are 95.22 and 87.78 respectively. This may be attributed to modeling multi-expert modal distribution and jointly discriminating misinformation news. Due to space limitations, the results of the Weibo21 dataset can be viewed in the appendix.
Answear to RQ2. Compared with previous methods of incomplete modality learning (Xu et al., 2024; Wu et al., 2024; Zeng et al., 2025), we first explored more fine-grained modality missing scenarios. Inspired by the previous methods, which explored two cases where the image and text modalities were 100% missing, we designed a total of 14 modality missing scenarios and conducted experiments, covering the previous modality missing settings. Specifically, we set the missing ratio $r$, where the text modality missing ratio is $r_{t}$ and the image modality missing ratio is $_{v}. r=r$$r_{t}+r_{v}$, where $r,r_{t},r_{v}\in\{0,25,50,75,100\}\subseteq[0,100]$. Here, we visualize the incompleteness of the five modalities when $r=100$, as shown in Figures 4, 5, and 6. Table 1 shows in detail the cases where text modality missing dominates ($r_{t}=75$, $r_{v}=25$) and image modality missing dominates ($r_{t}=25$, $r_{v}=75$) when $r=100$. Experiments show that MMLNet performs best in most cases of modality incompleteness and has better robustness. Analyzing the bad cases, in the weibo dataset, when the text modality missing is dominant, although the performance of MIMoE is slightly better than that of MMLNet, when the image modality missing is dominant, the performance of MIMoE deteriorates significantly, with the accuracy dropping by nearly 13%, while the drop of MMLNet is only 2.6%. By modeling the distribution of different modalities learned by different experts and learning the consistency of modality missing, MMLNet can cope with various modal missingness with ease. Please refer to the appendix for the complete experiment of modality missing scenario.
@@ 4.5. Main Results
Table 2. We conduct ablation experiments in two settings, incomplete and complete modalities, and the ablation of different modules verifies the effectiveness. In the incomplete modality setting, the text modality is missing 50% and the image modality is missing 50%. ‘w/o’ is anabbreviation for ‘with’, and ‘w/o’ is anabbreviation for ‘without’.
Method Complete Modality Incomplete Modality
Weibo Pheme Weibo Pheme
ACC(%) F1(%) AUC(%) ACC(%) F1(%) AUC(%) ACC(%) F1(%) AUC(%) ACC(%) F1(%) AUC(%)
NSLM 92.28 92.28 92.32 84.65 81.82 82.88 82.59 82.25 82.23 78.87 75.86 77.97
w/ ${L}_{m}$ 93.44 93.44 93.54 86.96 84.31 84.84 86.68 86.68 86.78 81.18 76.44 75.82
w/ AA 92.90 92.89 92.90 85.64 83.20 84.78 85.87 85.84 86.10 80.52 75.70 75.18
MIMOE 92.49 92.47 92.42 85.64 81.90 89.40 84.36 84.23 84.74 73.26 58.20 79.17
w/ ${L}_{m}$ 93.85 93.85 93.91 87.12 84.33 84.45 89.48 89.47 89.47 75.08 65.90 78.99
w/ AA 93.37 93.37 93.48 86.13 82.20 88.99 86.89 86.74 86.64 74.91 61.50 78.47
MMLNet-Ours 95.22 95.21 98.93 87.78 84.35 90.97 91.60 91.58 96.87 82.01 76.01 74.16
w/o ${L}_{c}^{h}$ 94.67 94.65 98.79 87.45 84.29 89.53 90.10 90.09 96.41 81.02 74.58 85.25
w/o ${L}_{c}^{r}$ 94.88 94.87 98.91 87.29 84.05 89.49 91.33 91.32 97.12 81.18 73.12 70.67
w/o ${L}_{c}^{f}$ 91.80 91.80 97.57 86.46 82.86 89.60 89.21 89.15 95.83 79.86 71.94 84.75
w/o ${L}_{m}^{h}$ 95.01 95.01 98.95 87.12 83.94 90.69 90.92 90.91 96.76 80.52 73.98 84.44
w/o ${L}_{m}^{r}$ 94.94 94.94 98.97 86.96 83.93 89.64 91.26 91.24 97.06 81.51 74.85 72.79
w/o ${L}_{m}^{f}$ 93.31 93.27 98.41 86.30 82.75 89.84 89.76 89.73 96.24 80.19 74.82 84.95
w/o AA 94.19 94.18 98.53 86.63 83.29 89.80 89.82 89.82 96.80 80.69 75.40 84.47
@@
Figure 7. Visualization of Ablation Experiments on Modality Missing Learning Module on the Weibo Dataset. Subfigure 7(a): Accuracy on the Testing Set. The Following Four Subfigures: Visualizations of T-SNE Dimensionality Reduction of Multimodal Features Before Classification Corresponding to the Four Curves in Subfigure 7(a).
Figure 8. Case study on fake news detection. Here, we perform analysis on the Pheme dataset with both complete and incomplete modalities. ‘T’ represents the text modality and ‘I’ represents the image modality. ‘Text modality missing predominates’ refers to T↓75%, I↓25%. ‘Image modality missing predominates’ refers to T↓25%, I↓75%.:
4.6. Ablation Study

### v2

@@ 3.3.1 Expert of Text-Modality
$F_{T}=f_{A}^{T}(h_{\text{{[CLS]}}})$ (3)
${y}^{h}=\text{softmax}(\boldsymbol{W}F_{T}+\boldsymbol{b})$ (4)
@@
3.3.2 Expert of Image-Modality
@@
$F_{I}=f_{A}^{I}({r}_{\text{{[CLS]}}})$ (6)
${y}^{r}=\text{softmax}(\boldsymbol{W}F_{I}+\boldsymbol{b})$ (7)
@@
Figure 3: Distribution of features in space for complete and incomplete modality.
3.3.3 Expert of Multi-Modality
Modeling the distribution between text modality and image modality is advantageous for capturing deep-level deceptive information. However, incomplete modality information can mislead the modeling of inter-modality distributions, thereby hindering the recognition of deceptive information. In this work, we employ a transformer encoder to model the joint expert distribution across multiple modalities.
@@ 3.3.3 Expert of Multi-Modality
$F=(H,R)=(h_{\text{[CLS]}},h_{1},\dots,h_{m},r_{\text{[CLS]}},r_{1},\dots,r_{n})$ (8)
@@
$\hat{F}=(\hat{h}_{\text{[CLS]}},\hat{h}_{1},\dots,\hat{h}_{m},\hat{r}_{\text{[CLS]}},\hat{r}_{1},\dots,\hat{r}_{n})=$
$\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d_{k}}}\right)V$ (9)
@@
$p_{t},p_{v}=\text{softmax}(\boldsymbol{W}(\hat{h}_{\text{[CLS]}},\hat{r}_{\text{[CLS]}})+b)$ (10)
$f=p_{t}\hat{h}_{\text{[CLS]}}+p_{v}\hat{r}_{\text{[CLS]}})$ (11)
$y^{f}=\text{softmax}(\boldsymbol{W}f+\boldsymbol{b})$ (12)
@@
3.3.4 Multi-Expert Reasoning Paradigm
@@
${y}^{o}=\sum_{M\in\left\{h,r,f\right\}}{\lambda}_{o}^{M}{y}^{M},{\lambda}_{o}\in\left[0,1\right]$ (13)
@@
3.4 Incomplete Modality Adapters
Mainstream MMR tasks typically employ models pre-trained on large-scale datasets as encoders, which may lead to overfitting of feature distributions, a problem that becomes more pronounced when modality information is missing. Inspired by clip_ada , we introduce lightweight feature adapters for single-modality image and text, enabling the learning of new modality features to compensate for the missing information while preserving the original distributions learned by the pre-trained models.
@@ 3.4 Incomplete Modality Adapters
$F_{I}=\alpha A^{I}({r}_{\text{{[CLS]}}})+(1-\alpha){r}_{\text{{[CLS]}}}=f_{A}^{I}({r}_{\text{{[CLS]}}})$ (14)
$F_{T}=\alpha A^{T}({h}_{\text{{[CLS]}}})+(1-\alpha){h}_{\text{{[CLS]}}}=f_{A}^{T}({h}_{\text{{[CLS]}}})$ (15)
@@
3.5 Modality Missing Learning
@@
To enhance the performance of MMLNet in MMR tasks under conditions of incomplete modality information, we propose a modality missing learning algorithm based on a vanilla multimodal contrastive learning framework. This simple yet effective approach aims to mitigate the negative effects of missing modality information by constraining the feature distributions across multiple modalities. Specifically, for each training sample $x$, we first define the corresponding positive sample set $\mathcal{S}_{p}$ and negative sample set $\mathcal{S}_{n}$ for it. Here, $\mathcal{S}_{p}$ contains samples with the same label as $x$, while $\mathcal{S}_{n}$ contains samples with the opposite label as $x$. As shown in Figure 3, we aim to learn a robust latent representation of $x$ by attracting $x$ to the representation of samples in $\mathcal{S}_{p}$ and repidating it away from the representation of samples in $\mathcal{S}_{n}$ as follows:
$\mathcal{L}_{m}=\frac{1}{\left|\mathcal{S}_{p}\right|}\sum_{p\in\mathcal{S}_{p}}-\log\frac{\exp\left(f(\mathbf{h})\cdot f\left(\mathbf{h}_{p}\right)/\tau\right)}{\sum_{n\in\mathcal{S}_{n}}\exp\left(f(\mathbf{h})\cdot f\left(\mathbf{h}_{n}\right)/\tau\right)}$ (16)
@@
However, in incomplete modal MMR task, it is suboptimal to weight all positive and negative samples equally. This alignment objective is not fully compatible with multimodal misinformation recognition with incomplete modalities. Based on the aforementioned perceptions, our model encounters two common but difficult problems caused by the lack of modality information: 1) distinguishing samples with similar semantic information but opposite labels due to incomplete modal information, and 2) narrowing the gap between samples with different semantic information but the same label due to incomplete modal information. To accomplish this, we introduce a label-aware adaptive weighting strategy. Specifically, we measure the semantic similarity of two samples by measuring the cosine similarity, each sample is re-weighted as follows:
$w_{c}=\left\{\begin{array}[]{l}1-cos(h_{c},h),\left(c\in\mathcal{S}_{p}\right)\\ 1+cos(h_{c},h),\left(c\in\mathcal{S}_{n}\right)\end{array}\right$. (19)
@@
$\hat{\mathcal{L}}_{m}=\frac{1}{\left|\mathcal{S}_{p}\right|}\sum_{p\in\mathcal{S}_{p}}-\log\frac{w_{p}\cdot\exp\left(f(\mathbf{h})\cdot f\left(\mathbf{h}_{p}\right)/\tau\right)}{\sum_{n\in\mathcal{S}_{n}}w_{n}\cdot\exp\left(f(\mathbf{h})\cdot f\left(\mathbf{h}_{n}\right)/\tau\right)}$ (20)
@@
In fact, modality missing learning is not only effective for multimodal features. We also apply it to text modality features $F_{T}$ and image modality features $F_{I}$ to further constrain the distribution of unimodal features and enhance the robustness of the MMLNet model in the incomplete modality MMR task.
3.6 Overall Optimization
@@
$\mathcal{L}=\sum_{M\in\{I,T,F\}}\lambda_{o}^{M}\left(\lambda_{c}\mathcal{L}_{c}^{M}+\lambda_{m}\mathcal{L}_{m}^{M}\right)$ (22)
@@
4 Experiments
In this section, we conduct relevant experiments to address several research questions. Through these experimental results, we systematically evaluate the effectiveness of MMLNet in tackling key challenges in incomplete modality misinformation recognition:
  • RQ1: Can MMLNet effectively enhance the performance of misinformation recognition?
  • RQ2: Is the proposed design robust for misinformation recognition under incomplete modalities?
  • RQ3: Do individual components contribute to improving misinformation recognition capability?
  • RQ4: Can modallity missing learning significantly improve incomplete modality MMR?
@@ 4 Experiments
4.1 Datasets
We evaluate the performance of MMLNet on three real-world datasets: Weibo weibo , Weibo21 weibo21 , and Pheme pheme , covering both Chinese and English. The Weibo dataset is the most widely used Chinese misinformation recognition dataset, containing fake and real news from Sina Weibo between 2012 and 2016, verified by authoritative sources such as Xinhua News Agency and Sina’s official debunking system. The Weibo21 dataset, released in 2021, is an updated version of the Weibo dataset that includes more recent Weibo news. Pheme is an English-language dataset consisting of tweets from Twitter, focusing on five breaking news events. Each event comprises a set of posts containing a substantial amount of textual and visual content, along with corresponding annotations. Three events are randomly selected, and their associated tweets are used for training, while tweets from the remaining events are utilized for testing. For comparison with baseline methods, we follow previous studies NSLMaaai24 , MOE_WWW25 , DGAmm24 , pheme_pro and use the same preprocessed dataset.
(a) T↓0%, I↓100%.
(b) T↓25%, I↓75%.
(c) T↓50%, I↓50%.
(d) T↓75%, I↓25%.
(e) T↓100%, I↓0%.
Figure 4: Experimental results of Weibo dataset. ‘T’ and ‘I’ respectively represent the text and image modalities.
(a) T↓0%, I↓100%.
(b) T↓25%, I↓75%.
(c) T↓50%, I↓50%.
(d) T↓75%, I↓25%.
(e) T↓100%, I↓0%.
Figure 5: Experimental results of Weibo21 dataset. ‘T’ and ‘I’ respectively represent the text and image modalities.
(a) T↓0%, I↓100%.
(b) T↓25%, I↓75%.
(c) T↓50%, I↓50%.
(d) T↓75%, I↓25%.
(e) T↓100%, I↓0%.
Figure 6: Experimental results of Pheme dataset. ‘T’ and ‘I’ respectively represent the text and image modalities.
4.2 Evaluation Metrics
The multimodal misinformation recognition problem is a binary classification task. Based on previous work NSLMaaai24 , fkaowl , MOE_WWW25 , we apply the accuracy score (ACC), the area under the receiver operating characteristic curve (AUC), and the macro-F1 score (M-F1) as evaluation metrics.
4.3 Baseline
We compare the two open-sourced sota MFDN methods including NSLM NSLMaaai24 and MIMoE MOE_WWW25 . We also fine-tunes the parameters of MLLM to show the ability of MLLM in MMR in complete and incomplete modal scenarios, specifically using the open-sourced MLLM Qwen-VL-7B [^1] https://github.com/QwenLM/Qwen-VL and LLaVA-v1.5-13B [^2] https://huggingface.co/liuhaotian/llava-v1.5-13b.
4.4 Implementation Details
To facilitate future research on incomplete modality MMR, for handling missing text modality, we set a random seed and pre-store the missing text. For missing image modality, we store the indices of the missing patches. Specifically, the image is resized to $224\times 224$ and decompose it into patches of size $32\times 32$. For the simplicity of the experiment, the training set and the test set have the same missing settings.
@@ 4.4 Implementation Details
Table 2: Ablation experiments on complete modalities on the Weibo and Pheme datasets. ‘w/o’ is anabbreviation for ‘with’, and ‘w/o’ is anabbreviation for ‘without’.
Method Complete Modality
Weibo Pheme
ACC(%) F1(%) AUC(%) ACC(%) F1(%) AUC(%)
NSLM 92.28 92.28 92.32 84.65 81.82 82.88
w/ ${L}_{m}$ 93.44↑1.16 93.44↑1.16 93.54↑1.22 86.96↑2.31 84.31↑2.49 84.84↑1.96
w/ AA 92.90↑0.62 92.89↑0.61 92.90↑0.58 85.64↑0.99 83.20↑1.38 84.78↑1.90
MIMOE 92.49 92.47 92.42 85.64 81.90 89.40
w/ ${L}_{m}$ 93.85↑1.36 93.85↑1.38 93.91↑1.49 87.12↑1.48 84.33↑2.43 84.45↓4.95
w/ AA 93.37↑0.88 93.37↑0.90 93.48↑1.06 86.13↑0.49 82.20↑0.30 88.99↓0.41
MMLNet-Ours 95.22 95.21 98.93 87.78 84.35 90.97
w/o ${L}_{c}^{h}$ 94.67↓0.55 94.65↓0.56 98.79↓0.14 87.45↓0.33 84.29↓0.06 89.53↓1.44
w/o ${L}_{c}^{r}$ 94.88↓0.34 94.87↓0.34 98.91↓0.02 87.29↓0.49 84.05↓0.30 89.49↓1.48
w/o ${L}_{c}^{f}$ 91.80↓3.42 91.80↓3.41 97.57↓1.36 86.46↓1.32 82.86↓1.49 89.60↓1.37
w/o ${L}_{m}^{h}$ 95.01↓0.21 95.01↓0.20 98.95↑0.02 87.12↓0.66 83.94↓0.41 90.69↓0.28
w/o ${L}_{m}^{r}$ 94.94↓0.28 94.94↓0.27 98.97↑0.04 86.96↓0.82 83.93↓0.42 89.64↓1.33
w/o ${L}_{m}^{f}$ 93.31↓1.91 93.27↓1.94 98.41↓0.52 86.30↓1.48 82.75↓1.60 89.84↓1.13
w/o AA 94.19↓1.03 94.18↓1.03 98.53↓0.40 86.63↓1.15 83.29↓1.06 89.80↓1.17
@@
Table 3: Ablation experiments on incomplete modalities on the Weibo and Pheme datasets. In the incomplete modality setting, the text modality is missing 50% and the image modality is missing 50%. ‘w/o’ is anabbreviation for ‘with’, and ‘w/o’ is anabbreviation for ‘without’.
Method Incomplete Modality
Weibo Pheme
ACC(%) F1(%) AUC(%) ACC(%) F1(%) AUC(%)
NSLM 82.59 82.25 82.23 78.87 75.86 77.97
w/ ${L}_{m}$ 86.68↑4.09 86.68↑4.43 86.78↑4.55 81.18↑2.31 76.44↑0.58 75.82↓2.15
w/ AA 85.87↑3.28 85.84↑3.59 86.10↑3.87 80.52↑1.65 75.70↓0.16 75.18↓2.79
MIMOE 84.36 84.23 84.74 73.26 58.20 79.17
w/ ${L}_{m}$ 89.48↑5.12 89.47↑5.24 89.47↑4.73 75.08↑1.82 65.90↑7.70 78.99↓0.18
w/ AA 86.89↑2.53 86.74↑2.51 86.64↑1.9 74.91↑1.65 61.50↑3.30 78.47↓0.70
MMLNet-Ours 91.60 91.58 96.87 82.01 76.01 74.16
w/o ${L}_{c}^{h}$ 90.10↓1.50 90.09↓1.49 96.41↓0.46 81.02↓0.99 74.58↓1.43 85.25↑11.09
w/o ${L}_{c}^{r}$ 91.33↓0.27 91.32↓0.26 97.12↑0.25 81.18↓0.83 73.12↓2.89 70.67↑3.49
w/o ${L}_{c}^{f}$ 89.21↓2.39 89.15↓2.43 95.83↓1.04 79.86↓2.15 71.94↓4.07 84.75↑10.59
w/o ${L}_{m}^{h}$ 90.92↓0.68 90.91↓0.67 96.76↓0.11 80.52↓1.49 73.98↓2.03 84.44↑10.28
w/o ${L}_{m}^{r}$ 91.26↓0.34 91.24↓0.34 97.06↑0.19 81.51↓0.50 74.85↓1.16 72.79↑1.37
w/o ${L}_{m}^{f}$ 89.76↓1.84 89.73↓1.85 96.24↓0.63 80.19↓1.82 74.82↓1.19 84.95↑10.97
w/o AA 89.82↓1.78 89.82↓1.76 96.80↓0.07 80.69↓1.32 75.40↓0.61 84.47↑10.31
@@
Figure 7: Visualization of Ablation Experiments on Modality Missing Learning Module on the Weibo Dataset. Subfigure 7(a): Accuracy on the Testing Set. The Following Four Subfigures: Visualizations of T-SNE Dimensionality Reduction of Multimodal Features Before Classification Corresponding to the Four Curves in Subfigure 7(a).
4.5 Main Results
Answear to RQ1. We conducted experiments using Chinese and English datasets in the complete modality scenario, as shown in Table 1. MMLNet significantly outperforms existing sota MMR methods and MLLM in three evaluation indicators. On the Chinese weibo dataset, the accuracy scores of the sota MMR method and MLLM are improved by 2.73% and 12.22% respectively. On the English pheme dataset, the accuracy scores of the sota MMR method and MLLM are improved by 2.14% and 30.68% respectively. These experimental results prove that MMLNet can more accurately identify misinformation in news. For example, the accuracy scores on Weibo and Pheme are 95.22 and 87.78 respectively. This may be attributed to modeling multi-expert modal distribution and jointly discriminating misinformation news. Due to space limitations, the results of the Weibo21 dataset can be viewed in the appendix.
Answear to RQ2. Compared with previous methods of incomplete modality learning MM24robustSenDistribution , MM24robustSenMOE , IMOL , we first explored more fine-grained modality missing scenarios. Inspired by the previous methods, which explored two cases where the image and text modalities were 100% missing, we designed a total of 14 modality missing scenarios and conducted experiments, covering the previous modality missing settings. Specifically, we set the missing ratio $r$, where the text modality missing ratio is $r_{t}$ and the image modality missing ratio is $_{v}. r=r$$r_{t}+r_{v}$, where $r,r_{t},r_{v}\in\{0,25,50,75,100\}\subseteq[0,100]$. Here, we visualize the incompleteness of the five modalities when $r=100$, as shown in Figures 4, 5, and 6. Table 1 shows in detail the cases where text modality missing dominates ($r_{t}=75$, $r_{v}=25$) and image modality missing dominates ($r_{t}=25$, $r_{v}=75$) when $r=100$. Experiments show that MMLNet performs best in most cases of modality incompleteness and has better robustness. Analyzing the bad cases, in the weibo dataset, when the text modality missing is dominant, although the performance of MIMoE is slightly better than that of MMLNet, when the image modality missing is dominant, the performance of MIMoE deteriorates significantly, with the accuracy dropping by nearly 13%, while the drop of MMLNet is only 2.6%. By modeling the distribution of different modalities learned by different experts and learning the consistency of modality missing, MMLNet can cope with various modal missingness with ease. Please refer to the appendix for the complete experiment of modality missing scenario.
Figure 8: Case study on misinformation recognition. Here, we perform analysis on the Pheme dataset with both complete and incomplete modalities. ‘T’ represents the text modality and ‘I’ represents the image modality. ‘Text modality missing predominates’ refers to T↓75%, I↓25%. ‘Image modality missing predominates’ refers to T↓25%, I↓75%.
4.6 Ablation Study